Two cars start from same place and at the same place but it move in opposite direction, after 8 hours they are 1200 km apart. If one vehicle move 10km/h more than other , what are speeds per hour?
Answer by Boreal(15235)

speed of 1 car is x, speed of other car is x+10
their combined speed is 2x+10 km/hr, and that equals 1200 km in 8 hours
so 8(2x+10)=1200, and the units are km on both sides (that must occur)
16x+80=1200
16x=1120
x=70 km/hr
x+10=80 km/hr
combined speed is 150 km/hr apart and that is 1200 km after 8 hours